The vector potential at the position defined by vector r in a uniform magnetic field B can be represented
1
Expert's answer
2019-04-15T09:46:03-0400

The vector potential A at the position defined by vector r in a uniform magnetic field B can be represented as


"A=\frac {1}{2}\times B\times r"